# Self-Consistent Renormalized Spin-Wave Theory of Magnetic and Topological Transitions in Two-Dimensional Honeycomb Ferromagnets

### v1

@@ III.1 Field-Tuned Magnetization and Phase Transitions
As a representative system, we consider the honeycomb ferromagnet CrI3, for which exchange parameters have been estimated in Ref. [Chen18]. Typical values are $J=2.01$ meV, $J^{\prime}=0.16$ meV, $D=0.31$ meV, and anisotropy $A=0.22$ meV, which corresponds to 1.9 T. Equivalently, in units of $J$ one has $J^{\prime}/J\approx 0.08$, $D/J\approx 0.15$ and $A/J\approx 0.11$. Based on these values, it is reasonable to assume that these interactions are on the order of $0.1J$. These parameters set the reference energy scale for the following discussion of magnetization curves and phase transitions. We first examine the cases where $J^{\prime}=0$ and take $D/J=0.1$ and $A/J=0.25$. Note that even in the absence of an external field, i.e., $h=0$, the presence of magnetic anisotropy induces a gap in the excitation energy, enabling the realization of intrinsic two-dimensional ferromagnetism. This can be clearly seen from the magnon energy near the $\Gamma$ point in Fig. 2(c).
@@
In contrast, the lower branches were found by scanning over trial magnetization values $m_{0}$ and accepting solutions where the output magnetization $m_{\mathrm{out}}$ matched $m_{0}$ within a predefined tolerance. This approach reveals the full set of fixed points, including those not reachable via iterative self-consistency. Within the coexistence region, the lower branch shows an unphysical increase of $m$ with temperature, signaling its thermodynamic instability (see Appendix B for further discussion). As a consequence, the system exhibits a discontinuous drop in magnetization at the bifurcation point ($T/J\approx 2.72$ for $h=0$), characteristic of a first-order transition. Above this temperature, only the trivial paramagnetic solution $m=0$ remains, whereas below $T_{\rm Curie}$ both stable and unstable solutions coexist within a certain temperature range. This behavior underscores the limitations of SRSWT, which predicts a first-order magnetic phase transition, even though a continuous (second-order) transition is typically expected for intrinsic two-dimensional ferromagnets [Chen20].
@@
As discussed in Sec. II, the topological phase transition temperature occurs when $m=m_{c}$. For $S=3/2$, we previously found $m_{c}=3/26\approx 0.115$ when the six-operator interaction term $\mathcal{H}_{6}$ is neglected [Lu24]. We now provide an improved estimate for the critical magnetization at which the gap closes, based on the full SRSWT framework that incorporates $\mathcal{H}_{6}$. The Dirac gap is solely determined from
@@ III.2 Spectral Gap and Stability of the Magnetic Phase
To further investigate the thermodynamic behavior and its connection to magnon excitations, we analyze the spectral gap at the $\Gamma$ point and its correlation with the stability of the magnetic phase. As seen in Fig. 2(a), the coexistence region—where multiple magnetization solutions exist—shrinks as the external field $h$ decreases. This trend correlates with the $\Gamma$-point spectral gap $\Delta_{\Gamma}$, which reflects the system’s resistance to long-wavelength fluctuations. In Fig. 3(a), we plot $\Delta_{\Gamma}$ as a function of temperature for the three external fields in Fig. 2(a), keeping $A/J=0.25$ fixed. The gap decreases as $h$ is reduced, suggesting that a smaller $\Delta_{\Gamma}$ is associated with a smaller coexistence region. At low temperature, where corrections from $\mathcal{H}_{4}$ and $\mathcal{H}_{6}$ are negligible, the spectral gap approaches the free-theory value $\Delta_{\Gamma}=2A+h$ for $S=3/2$, consistent with our results. Interestingly, $\Delta_{\Gamma}$ exhibits a strong non-monotonic dependence on temperature and typically reaches a minimum just below $T_{\rm Curie}$. While reducing $h$ can shrink the unstable region by suppressing the gap, overly negative $h$ can drive $\Delta_{\Gamma}\to 0$ at a finite temperature, rendering the theory invalid due to the onset of strong infrared fluctuations—akin to the mechanism described by the Mermin–Wagner theorem.
@@
Finally, in Fig. 3(c), we plot the corresponding $\Delta_{\Gamma}$ curves for the parameter sets in Fig. 3(b). For $A=0.2$, the minimum in $\Delta_{\Gamma}$ occurs at lower temperatures compared to $A=0.3$. Furthermore, in the $A=0.3$ case, the gap decreases monotonically with temperature, suggesting a smoother, more conventional thermal suppression of order. Taken together, these results indicate that for sufficiently large anisotropy, SRSWT yields behavior more consistent with a continuous magnetic phase transition, even though the transition remains technically first-order within this framework.
@@ III.3 Role of $J^{\prime}$ and DMI in Modifying the Magnetic Transition
In Fig. 4(b), we examine the same scenario with a finite Zeeman field $h/J=0.1$. Again, the repeated $J^{\prime}=0$ case is shown for reference. Remarkably, introducing a modest antiferromagnetic NNN interaction ($J^{\prime}/J=-0.07$) nearly eliminates the coexistence region, resulting in an almost continuous phase transition. This negative $J^{\prime}$ is not merely a model-specific choice. Comparable antiferromagnetic third-nearest-neighbor couplings have been extracted from inelastic neutron scattering in CrI3 [Chen21] and supported by DFT calculations [Cong22], lending experimental plausibility to this tuning regime. Our results illustrate that a suitably chosen negative $J^{\prime}$ can counteract the stabilizing effect of the external field and suppress the first-order nature of the transition. Physically, this can be attributed to the frustration introduced by negative NNN exchange, which weakens the long-range magnetic order and flattens the free energy landscape near the critical point. As a result, the phase transition becomes less abrupt and behaves as a continuous transition.
@@ IV Conclusion
A central result of our study is that, under suitable tuning—including an optimal combination of anisotropy and Zeeman field, along with the introduction of an antiferromagnetic next-nearest-neighbor (NNN) interaction—topological magnon band inversions driven by thermal fluctuations can occur within the stable, upper branch of the magnetization curve. This contrasts with the untuned case, where band gap closings appear only in metastable regimes and are thus not physically realizable. We explicitly demonstrated this for a case with negative $J^{\prime}$, where the Dirac-point magnon gap closes at a temperature lower than the Curie temperature, signaling a topological transition that is experimentally accessible. This negative $J^{\prime}$ is not merely a model-specific choice. In Ref. [Cong22], it was numerically found that the third-nearest-neighbor exchange can be antiferromagnetic. We further showed that DMI plays a subtle role: while it leaves the bulk magnetization nearly unchanged, it can enhance fluctuation effects sufficiently to convert an almost continuous transition into a discernibly weak first-order one.
@@
To assess the experimental feasibility of our predictions, we compare the characteristic temperature scales obtained from our theory with those reported in experiments on CrI3 [Chen18]. From Fig. 4(b), we find that the Curie temperature is approximately $T_{\rm Curie}\approx 2.5J$ for $J^{\prime}/J=-0.07$. Using $J=2.01$ meV from Ref. [Li18], this corresponds to $T_{\rm Curie}\simeq 58$ K, which is close to the experimental value $61$ K reported in the same work. In this parameter regime we also obtain a topological transition at $T_{c}\approx 2.4J$, i.e., about 2.3 K below $T_{\rm Curie}$. We chose this case because it yields the largest separation between $T_{c}$ and $T_{\rm Curie}$ within our parameter scan, making it the most favorable scenario for experimental detection. Such a difference should be experimentally accessible, for example through thermal Hall measurements [Onose10, Hirschberger15] or inelastic neutron scattering [Chen18].
@@ Appendix A Higher-order interacting Hamiltonian
The first thing we notice is that all of these interaction terms are number-conserving. This conservation follows from the global $U(1)$ spin-rotation symmetry about the $z$ axis, $\hat{S}^{+}\to e^{i\theta}\hat{S}^{+}$ and $\hat{S}^{-}\to e^{-i\theta}\hat{S}^{-}$, which enforces the conservation of the total $\hat{S}^{z}$ and, in the bosonic representation, the magnon number. As a direct consequence of this symmetry, anomalous thermal averages such as $\langle aa\rangle$ or $\langle a^{\dagger}a^{\dagger}\rangle$ vanish identically. Secondly, the terms with the prefactor $\frac{1}{32S^{2}}$ in $\mathcal{H}_{4}$ are actually derived from the reduction of the six-operator interaction in the process of normal ordering. These terms have important implications when comparing $T_{c}$ with $T_{\rm Curie}$ [Lu24]; see the discussion in Sec. II for details.
@@
$\approx$ $\sum_{{\bf k}}\left(12\rho_{{\bf p}}+6\rho_{{\bf k}}\right)\left\langle a_{{\bf p}}^{\dagger}a_{{\bf p}}\right\rangle\left\langle a_{{\bf q}}^{\dagger}a_{{\bf q}}\right\rangle a_{{\bf k}}^{\dagger}a_{{\bf k}}.$ (10)

### v2

@@ III.1 Field-Tuned Magnetization and Phase Transitions
As a representative system, we consider the honeycomb ferromagnet CrI3, for which exchange parameters have been estimated in Ref. [7]. Typical values are $J=2.01$ meV, $J^{\prime}=0.16$ meV, $D=0.31$ meV, and anisotropy $A=0.22$ meV, which corresponds to 1.9 T. Equivalently, in units of $J$ one has $J^{\prime}/J\approx 0.08$, $D/J\approx 0.15$ and $A/J\approx 0.11$. Based on these values, it is reasonable to assume that these interactions are on the order of $0.1J$. These parameters set the reference energy scale for the following discussion of magnetization curves and phase transitions. We first examine the cases where $J^{\prime}=0$ and take $D/J=0.1$ and $A/J=0.25$. Note that even in the absence of an external field, i.e., $h=0$, the presence of magnetic anisotropy induces a gap in the excitation energy, enabling the realization of intrinsic two-dimensional ferromagnetism. This can be clearly seen from the magnon energy near the $\Gamma$ point in Fig. 2(c).
@@
In contrast, the lower branches were found by scanning over trial magnetization values $m_{0}$ and accepting solutions where the output magnetization $m_{\mathrm{out}}$ matched $m_{0}$ within a predefined tolerance. This approach reveals the full set of fixed points, including those not reachable via iterative self-consistency. Within the coexistence region, the lower branch shows an unphysical increase of $m$ with temperature, signaling its thermodynamic instability (see Appendix B for further discussion). As a consequence, the system exhibits a discontinuous drop in magnetization at the bifurcation point ($T/J\approx 2.72$ for $h=0$), characteristic of a first-order transition. Above this temperature, only the trivial paramagnetic solution $m=0$ remains, whereas below $T_{\rm Curie}$ both stable and unstable solutions coexist within a certain temperature range. This behavior underscores the limitations of SRSWT, which predicts a first-order magnetic phase transition, even though a continuous (second-order) transition is typically expected for intrinsic two-dimensional ferromagnets [51].
@@
As discussed in Sec. II, the topological phase transition temperature occurs when $m=m_{c}$. For $S=3/2$, we previously found $m_{c}=3/26\approx 0.115$ when the six-operator interaction term $\mathcal{H}_{6}$ is neglected [48]. We now provide an improved estimate for the critical magnetization at which the gap closes, based on the full SRSWT framework that incorporates $\mathcal{H}_{6}$. The Dirac gap is solely determined from
@@ III.2 Spectral Gap and Stability of the Magnetic Phase
To further investigate the thermodynamic behavior and its connection to magnon excitations, we analyze the spectral gap at the $\Gamma$ point and its correlation with the stability of the magnetic phase. As seen in Fig. 2(a), the coexistence region—where multiple magnetization solutions exist—shrinks as the external field $h$ decreases. This trend correlates with the $\Gamma$-point spectral gap $\Delta_{\Gamma}$, which reflects the system’s resistance to long-wavelength fluctuations. In Fig. 3(a), we plot $\Delta_{\Gamma}$ as a function of temperature for the three external fields in Fig. 2(a), keeping $A/J=0.25$ fixed. The gap decreases as $h$ is reduced, suggesting that a smaller $\Delta_{\Gamma}$ is associated with a smaller coexistence region. At low temperature, where corrections from $\mathcal{H}_{4}$ and $\mathcal{H}_{6}$ are negligible, the spectral gap approaches the free-theory value $\Delta_{\Gamma}=2A+h$ for $S=3/2$, consistent with our results. Interestingly, $\Delta_{\Gamma}$ exhibits a strong non-monotonic dependence on temperature and typically reaches a minimum just below $T_{\rm Curie}$. The origin of this non-monotonic behavior is discussed in Appendix C. While reducing $h$ can shrink the unstable region by suppressing the gap, overly negative $h$ can drive $\Delta_{\Gamma}\to 0$ at a finite temperature, rendering the theory invalid due to the onset of strong infrared fluctuations—akin to the mechanism described by the Mermin–Wagner theorem.
@@
Next, in Fig. 3(c), we plot the corresponding $\Delta_{\Gamma}$ curves for the parameter sets in Fig. 3(b). For $A=0.2$, the minimum in $\Delta_{\Gamma}$ occurs at lower temperatures compared to $A=0.3$. Furthermore, in the $A=0.3$ case, the gap decreases monotonically with temperature, suggesting a smoother, more conventional thermal suppression of order. Taken together, these results indicate that for sufficiently large anisotropy, SRSWT yields behavior more consistent with a continuous magnetic phase transition, even though the transition remains technically first-order within this framework.
The selection of the external Zeeman field $h$ is governed by the self-consistent interplay between $\Delta_{\Gamma}$ and the uniform magnetization, which ultimately determines the accessibility of the topological phase transition. For moderate anisotropy ($A=0.25$), the use of a negative opposing field (h=0.07h=-0.07) is essential. Physically, the negative field reduces the energy cost of fluctuations at the zone center, leading to a smaller $\Delta_{\Gamma}$. In our self-consistent framework, this reduction in $\Delta_{\Gamma}$ enhances the thermal population of magnons, thereby driving the magnetization $m$ to smaller values. This suppression of $m$ is a critical prerequisite for the topological transition; the renormalization effects required to close the Dirac gap $\Delta_{K}$ can only be achieved when the magnetization is sufficiently reduced (recall that $m_{c}\approx 0.238$). If $h$ were zero or positive in this regime, $\Delta_{\Gamma}$ would remain large, keeping $m$ too high to access the gap-closing point on the physical solution branch. However, $h$ cannot be made arbitrarily negative, as an excessive reduction in $\Delta_{\Gamma}$ would cause the stability gap to collapse before the topological gap closes as mentioned earlier in this subsection. Thus, h=0.07h=-0.07 represents the optimal balance for $A=0.25$, ensuring $m$ is small enough to permit the transition but finite enough to maintain order.
Conversely, when the single-ion anisotropy is increased to $A=0.3$, the stability dynamics are inverted. Although anisotropy is typically associated with stabilizing the magnetic order, in the self-consistent framework, a larger $A$ significantly enhances the softening effects which leads to a premature collapse of the stability gap ($\Delta_{\Gamma}$) at a relatively high magnetization value (see Appendix C), thereby destabilizing the ferromagnetic state well before the topological transition can occur. In this regime, a positive Zeeman field ($h=0.1$) becomes the preferred configuration. Unlike the $A=0.25$ case where a negative field was needed to soften the spectrum, here the positive field is required to counteract the anisotropy-driven instability. By providing a rigid energy offset that opposes the rapid renormalization of $\Delta_{\Gamma}$, the positive field prevents the premature gap closing. This stabilizing effect successfully opens a narrow, yet physical, temperature window where the topological transition is observable prior to the thermal destruction of the ferromagnetic state.
@@ III.3 Role of $J^{\prime}$ and DMI in Modifying the Magnetic Transition
In Fig. 4(b), we examine the same scenario with a finite Zeeman field $h/J=0.1$. Again, the repeated $J^{\prime}=0$ case is shown for reference. Remarkably, introducing a modest antiferromagnetic NNN interaction ($J^{\prime}/J=-0.07$) nearly eliminates the coexistence region, resulting in an almost continuous phase transition. This negative $J^{\prime}$ is not merely a model-specific choice. Comparable antiferromagnetic third-nearest-neighbor couplings have been extracted from inelastic neutron scattering in CrI3 [8] and supported by DFT calculations [52], lending experimental plausibility to this tuning regime. Our results illustrate that a suitably chosen negative $J^{\prime}$ can counteract the stabilizing effect of the external field and suppress the first-order nature of the transition. Physically, this can be attributed to the frustration introduced by negative NNN exchange, which weakens the long-range magnetic order and flattens the free energy landscape near the critical point. As a result, the phase transition becomes less abrupt and behaves as a continuous transition.
@@ IV Conclusion
A central result of our study is that, under suitable tuning—including an optimal combination of anisotropy and Zeeman field, along with the introduction of an antiferromagnetic next-nearest-neighbor (NNN) interaction—topological magnon band inversions driven by thermal fluctuations can occur within the stable, upper branch of the magnetization curve. This contrasts with the untuned case, where band gap closings appear only in metastable regimes and are thus not physically realizable. We explicitly demonstrated this for a case with negative $J^{\prime}$, where the Dirac-point magnon gap closes at a temperature lower than the Curie temperature, signaling a topological transition that is experimentally accessible. This negative $J^{\prime}$ is not merely a model-specific choice. In Ref. [52], it was numerically found that the third-nearest-neighbor exchange can be antiferromagnetic. We further showed that DMI plays a subtle role: while it leaves the bulk magnetization nearly unchanged, it can enhance fluctuation effects sufficiently to convert an almost continuous transition into a discernibly weak first-order one.
@@
To assess the experimental feasibility of our predictions, we compare the characteristic temperature scales obtained from our theory with those reported in experiments on CrI3 [7]. From Fig. 4(b), we find that the Curie temperature is approximately $T_{\rm Curie}\approx 2.5J$ for $J^{\prime}/J=-0.07$. Using $J=2.01$ meV from Ref. [40], this corresponds to $T_{\rm Curie}\simeq 58$ K, which is close to the experimental value $61$ K reported in the same work. In this parameter regime we also obtain a topological transition at $T_{c}\approx 2.4J$, i.e., about 2.3 K below $T_{\rm Curie}$. We chose this case because it yields the largest separation between $T_{c}$ and $T_{\rm Curie}$ within our parameter scan, making it the most favorable scenario for experimental detection. Such a difference should be experimentally accessible, for example through thermal Hall measurements [53, 32] or inelastic neutron scattering [7]. This large separation explicitly demonstrates that the topological gap closing is a distinct physical mechanism driven by band renormalization, independent of the critical fluctuations associated with $T_{\rm Curie}$. Consequently, the proximity of $T_{\text{Curie}}$ and $T_{c}$ observed in the $J^{\prime}>0$ case (Fig. 2) should be understood as a specific characteristic of that parameter space, rather than an intrinsic coupling of the two transitions.
@@ Appendix A Higher-order interacting Hamiltonian
The first thing we notice is that all of these interaction terms are number-conserving. This conservation follows from the global $U(1)$ spin-rotation symmetry about the $z$ axis, $\hat{S}^{+}\to e^{i\theta}\hat{S}^{+}$ and $\hat{S}^{-}\to e^{-i\theta}\hat{S}^{-}$, which enforces the conservation of the total $\hat{S}^{z}$ and, in the bosonic representation, the magnon number. As a direct consequence of this symmetry, anomalous thermal averages such as $\langle aa\rangle$ or $\langle a^{\dagger}a^{\dagger}\rangle$ vanish identically. Secondly, the terms with the prefactor $\frac{1}{32S^{2}}$ in $\mathcal{H}_{4}$ are actually derived from the reduction of the six-operator interaction in the process of normal ordering. These terms have important implications when comparing $T_{c}$ with $T_{\rm Curie}$ [48]; see the discussion in Sec. II for details.
@@
$\approx$ $\sum_{\bf k,p,q}\left(12\rho_{{\bf p}}+6\rho_{{\bf k}}\right)\left\langle a_{{\bf p}}^{\dagger}a_{{\bf p}}\right\rangle\left\langle a_{{\bf q}}^{\dagger}a_{{\bf q}}\right\rangle a_{{\bf k}}^{\dagger}a_{{\bf k}}.$ (10)
@@ Appendix B Free energy
Appendix C Decomposition of Self-Energy Contributions at the $\Gamma$ Point
To elucidate the physical origin of the anomalous magnon hardening observed at the $\Gamma$ point ($\mathbf{k}=0$), we perform a decomposition of the self-consistent energies. In the SRSWT formalism, the temperature-dependent gap $\Delta(T)$ can be expressed as the sum of the temperature-independent bare gap $E^{(0)}$ (determined by the Zeeman field and anisotropy) and the renormalization corrections arising from the Hartree-Fock decoupling of the quartic ($H_{4}$) and sextic ($H_{6}$) interaction terms, denoted as $\delta E^{(4)}(T)$ and $\delta E^{(6)}(T)$, respectively. As illustrated in Fig. 6, these two contributions exhibit competing behaviors as thermal fluctuations increase. The quartic correction $\delta E^{(4)}(T)$ is dominated by the single-ion anisotropy and provides a negative contribution to the energy. Consistent with the extensive discussion in Ref. [17], we find that truncating the theory at this order leads to a spurious, rapid softening of the magnon gap. Specifically, if only $H_{4}$ corrections are retained, the gap is predicted to close prematurely at a magnetization of $m\approx 1$ (for $S=3/2$), implying a phase transition at a temperature where the system is still highly ordered, which is physically unlikely. However, the inclusion of the sextic term $\delta E^{(6)}(T)$ introduces a positive energy shift that grows with the thermal magnon population. As temperature increases, this positive contribution from $H_{6}$ effectively competes with and eventually dominates the anisotropy-driven softening from $H_{4}$. Consequently, the net effect is a hardening of the $\Gamma$-point energy, which prevents the unphysical early gap closing and stabilizes the ferromagnetic order up to the true critical temperature. Thus, the observed hardening is not a numerical artifact, but a necessary regularization provided by higher-order interactions to maintain thermodynamic stability in the presence of strong single-ion anisotropy.
Figure 6: We use the same parameters as those of the red curve in Fig. 3(c). Shown is the decomposition of the magnon energy at the $\Gamma$ point as a function of temperature. The red curve represents the energy including only the bare gap and quartic corrections, $E^{(0)}+\delta E^{(4)}$, illustrating a spurious softening that would lead to an unphysical instability (a negative gap) at $T/J\approx 1.5$. The blue curve shows the contribution from the sextic corrections, $\delta E^{(6)}$, which grows with increasing temperature. Note that the dominance of $\delta E^{(6)}$ over the red curve at higher temperatures is the necessary physical condition for thermodynamic stability, analogous to the stabilizing role of positive higher-order terms in a Ginzburg-Landau expansion. The green curve corresponds to the fully renormalized energy, $E^{(0)}+\delta E^{(4)}+\delta E^{(6)}$. The observed hardening at higher temperatures arises from the dominant stabilizing effect of $\delta E^{(6)}$, which rescues the system from the artificial instability induced by the truncated expansion.